solve the trigonometric equation tanA+tan2A+(3^1/2)tanAtan2=3^1/2 

Grade:11

1 Answers

BALA SRIVATSAV YERRAMILLI
3 years ago
Heyy! Here’s your answer.
tanA + tan2A = √3 (1 – tanAtan2A),
(tanA + tan2A) / (1 – tanAtan2A) = √3,
tan(A + 2A) = tan π/3,
3A = nπ + π/3,
Therefore, A = nπ/3 + π/9.
Hope this helps you!